1. Alex Bowen has borrowed $40,000 to be repaid over 20 years at an interest rate of 9% p.a. What will be the annual installments to repay this loan?

2. The Pinewood Company is trying to determine the optimal order quantity of sofas. Annual sales of sofas are 900 units at a retail price of $300. The cost of carrying sofas is $1 per month per sofa. It costs $37.50 to prepare and receive an order. Determine the economic order quantity.

3. A company sells 60,000 litres of a commodity per year. Cost of the commodity is $10 per litre. Inventory carrying costs are 15% of the commodity cost. The cost of ordering is $72 per order. Order lead time averages five days and average daily demand is 200 liters per business day. Management wants a safety stock of ten days. Find reorder point.
Answer #1
1) The annual installments using the loan amortization formula = 40000*0.09*1.09^20/(1.09^20-1) = $ 4,381.86
2) EOQ is given by the formula:
EOQ = (2*A*O/c)^0.5, where
A = Annual sales, O = Ordering costs per order
and c = dollar cost of carrying 1 unit per year.
Substituting available values,
EOQ = (2*900*37.5/12)^0.5 = 75 Sofas
3) Re-order point = Safety stock+Lead time consumption
= 200*5+200*10 = 3000 litres